Given that z = ax +sqrt (a^2 − 4y) + c is the complete integral of the PDE, p^2 − q^2 =4
determine its general integral.
1
Expert's answer
2020-05-21T12:21:57-0400

Correction.

To understand the recall some theory.

Equation of the form "f(p,q)=0" i.e. the equation contains "p" and "q" only.

Suppose that "z(x,y)=ax+by+c" is solution of equation, then



"p=\frac{\partial z}{\partial x}=a\,\,\,\text{and}\,\,\,q=\frac{\partial z}{\partial y}=b"

substitute the above in equation "f(p,q)=0" , we get "f(a,b)=0" on solving this we can get "b=\phi(a)" , where "\phi" is known function. Using "b=\phi(a)" , the complete solution of the given partial equation is



"z=ax+\phi(a)y+c"

In our case,



"p^2-q^2=4\longrightarrow\underbrace{ p^2-q^2-4}_{f(p,q)}=0"

Then, suppose



"z=ax+by+c\\[0.3cm]
p=\frac{\partial z}{\partial x}=a\,\,\,\text{and}\,\,\,q=\frac{\partial z}{\partial y}=b\\[0.3cm]
p^2-q^2-4=0\longrightarrow a^2-b^2-4=0\longrightarrow\\[0.3cm]
\boxed{b=\pm\sqrt{a^2-4},\,\,\,\forall |a|\ge2}"

Conclusion,



"\boxed{z=ax\pm\sqrt{a^2-4}\cdot y+c\,\,\,-\,\,\text{is the general integral}}"



and



"z=ax+\sqrt{a^2-4}\cdot y+c"

is the complete integral as a special case.